Bear Normal Bull Market Market Market Probability 15.00% 50.00% 35.00% Stock X -12.00% 10.00% 21.00% Stock...

Bear

Normal

Bull

Market

Market

Market

Probability

15.00%

50.00%

35.00%

Stock X

-12.00%

10.00%

21.00%

Stock Y

-22.00%

14.00%

39.00%

3.c) If the risk–free rate of return is 2.95%, what are the Sharpe Ratios for stocks X and Y (8 points)? (Please assume that the standard deviations of the excess returns are the same as the standard deviations of returns calculated in part b)

Given that-
i ii iii iv=i*ii v=i*iii vi=i*(ii-10.55%%)^2 vii=i*(iii-17.35%)^2
Probability X Y Expected return X Expected return Y Variance A Variance B
15% -12% -22% -1.80% -3.30% 0.76% 2.32%
50% 10% 14% 5.00% 7.00% 0.00% 0.06%
35% 21% 39% 7.35% 13.65% 0.38% 1.64%
10.55% 17.35% 1.15% 4.02%
Expected return X 10.55%
Expected return Y 17.35%
Answer b. Expected SD of X = 1.15%^0.5 10.71%
Expected SD of Y = 4.02%^0.5 20.05%

Comment
Answer #1

SHARP RATIO FOR STOCK X = (10.55 - 2.95)/10.71 = 0.70

SHARP RATIO FOR STOCK Y = (17.35 - 2.95)/20.05 = 0.71